Status Report April 10, 2002 (LMG02-03)

From LMG Chief Scientist, Jose Torres to NBP Chief Scientist, Peter Wiebe

 

Hi Pete,

 

Ready for more good news? Things are still a wee bit bumpy here. The winds have been out of the SW for most of the day with speeds oscillating between 30 and 45 knots, mostly in the 30-35 range. Seas 15-20. We have been jitterbugging our way through the day. Barometer is holding steady at about 985. Our position now is 61S 61W and we are over the convergence, methinks, with SST about 0.3.

 

I will be having a PI meeting tomorrow about sampling locations in Crystal Sound and will let you know what we decide. The most likely locations are where the seals were doing the most hunting last trip, which appeared to be in the troughs vic Laird Island. I have attached a new straw plan more in keeping with our jointly worked out plan but taking into account the loss of a day on the front end.

 

Things seem to be calming down a bit right now. Winds have died down to 25. More tomorrow.

 

Best to all, Jose